OpenStudy (anonymous):

Janise rolled a number cube 50 times and recorded the results in this table. Outcome Number 1 7 2 9 3 8 4 8 5 6 6 12 What is the theoretical probability of the cube landing on a 6? What is the experimental probability of the cube landing on a 6?

OpenStudy (anonymous):

Theoretical probability is 1/6 (= approx. 0.17) Experimental probability (from the given table) is 12/50=6/25 (= approx. 0.24)

OpenStudy (anonymous):

Thanks! @HappySoul

OpenStudy (anonymous):

u r welcome